OpenStudy (anonymous):

What is the equation of the line passing through (8, –6) and perpendicular to the line y = x – 2 in slope-intercept form? a. y = –x + 2 b. y = x + 2 c. y = –x – 2 d. y = x – 2

OpenStudy (anonymous):

@Noura11 can you help?

OpenStudy (anonymous):

@shamim how about you?

OpenStudy (shamim):

f9

OpenStudy (shamim):

ur given line is\[y=x-2\]

OpenStudy (anonymous):

An equation of line is : \[y=mx+b\] Since the line is perpendicular to the line y = x – 2 then : \[m\times 1=-1\] and so : \[m=-1\] to find be, since (8,-6) is from the line : \[-6=-1\times 8+b\] so : \[b=-6+8=-2\] So the equation of the line is : \[y=-x-2\]

OpenStudy (anonymous):

Damn, that was amazing.

OpenStudy (shamim):

so the slope of ur given line is\[m _{1}=1\]

OpenStudy (anonymous):

1 problem. that answer wasnt one of the choices

OpenStudy (anonymous):

@Noura11

OpenStudy (anonymous):

What it is ?

OpenStudy (anonymous):

lol pellet, i was confusing 2 different problems, you are both correct. best answer goes to @Noura11 for the most correct work shown.